OpenStudy (anonymous):

Hello everyone, I have a question. I am supposed to put this into standard form, (2+i)/(3i). And to do so I have to multiply by the conjugate, how do i find the conjugate?

OpenStudy (anonymous):

Multiply by (2-i)/(2-i)

OpenStudy (anonymous):

Both the top and the bottom?

OpenStudy (anonymous):

Don't overthink it. Just multiply it.

myininaya (myininaya):

\[\frac{2+i}{3i}=\frac{2}{3i}+\frac{i}{3i}=\frac{2}{3i}*\frac{i}{i}+\frac{1}{3}=\frac{2i}{3i^2}+\frac{1}{3}\] \[\frac{2i}{3(-1)}+\frac{1}{3}\]

myininaya (myininaya):

\[=\frac{-2}{3}i+\frac{1}{3}\]

OpenStudy (anonymous):

Thank you. I made a mistake in my computation and you helped me catch it. Thanks :)

myininaya (myininaya):

np